OpenStudy (anonymous):

Let f(x) = x5 + x + 7. Find the the value of the inverse function at the given point. Someone please show me how to do this?

OpenStudy (anonymous):

what is the given point?

OpenStudy (anonymous):

oh 41

OpenStudy (anonymous):

f^-1 (41)

OpenStudy (nikvist):

\[f(x) = x^5 + x + 7\]\[f(2) = 2^5 + 2 + 7=41\quad\Rightarrow\quad f^{-1}(41)=2\]